OpenStudy (anonymous):

with head wind, a plane flew 3600 km in 10 hr the return flight over the same route took only 9hr. find the wind speed and planes air speed

OpenStudy (anonymous):

Let the plane speed be x Let the wind speed be b 10 (x - b) = 3600 x - b = 360 9 ( x + b) = 3600 x + b = 400 add the two equations 2x = 760 x = 380 b = 20

OpenStudy (anonymous):

plane speed is 380km / hr wind speed is 20 km/ hr